OpenStudy (anonymous):

Correct me . ? Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = 5 over 3; y-intercept = (0, 2) <----- slope = -5; y-intercept = (0, -6) slope = negative 3 over 5; y-intercept = (0, 2) slope = 5 over 3; y-intercept = (0, -6)

OpenStudy (anonymous):

why I choose the answer

OpenStudy (anonymous):

\[3y=5x+6....y=\frac{ 5 }{3}x +\frac{ 6 }{ 3 }......y=\frac{ 5 }{3}x +2\]